# Hate Speech Detection with Generalizable Target-aware Fairness

### v1

@@ Abstract.
To counter the side effect brought by the proliferation of social media platforms, hate speech detection (HSD) plays a vital role in halting the dissemination of toxic online posts at an early stage. However, given the ubiquitous topical communities on social media, a trained HSD classifier can easily become biased towards specific targeted groups (e.g., female and black people), where a high rate of either false positive or false negative results can significantly impair public trust in the fairness of content moderation mechanisms, and eventually harm the diversity of online society. Although existing fairness-aware HSD methods can smooth out some discrepancies across targeted groups, they are mostly specific to a narrow selection of targets that are assumed to be known and fixed. This inevitably prevents those methods from generalizing to real-world use cases where new targeted groups constantly emerge (e.g., new forums created on Reddit) over time. To tackle the defects of existing HSD practices, we propose Generalizable target-aware Fairness (GetFair), a new method for fairly classifying each post that contains diverse and even unseen targets during inference. To remove the HSD classifier’s spurious dependence on target-related features, GetFair trains a series of filter functions in an adversarial pipeline, so as to deceive the discriminator that recovers the targeted group from filtered post embeddings. To maintain scalability and generalizability, we innovatively parameterize all filter functions via a hypernetwork. Taking a target’s pretrained word embedding as input, the hypernetwork generates the weights used by each target-specific filter on-the-fly without storing dedicated filter parameters. In addition, a novel semantic gap alignment scheme is imposed on the generation process, such that the produced filter function for an unseen target is rectified by its semantic affinity with existing targets used for training. Finally, experiments are conducted on two benchmark HSD datasets, showing advantageous performance of GetFair on out-of-sample targets among baselines.
@@
journalyear: 2024copyright: acmcopyrightconference: Proceedings of the 27th ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 25–29, 2024; Barcelona, Spain.booktitle: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining (KDD ’24), August 25–29, 2024, Barcelona, Spainprice: 15.00copyright: noneccs: Information systems Web miningccs: Information systems Data miningccs: Security and privacy Social aspects of security and privacy
@@ 1. Introduction
To this end, in this paper we propose a novel approach, namely HSD with Generalizable target-aware Fairness (GetFair). In GetFair, to avoid the linear parameter cost associated with the number of target-specific filter functions, instead of independently parameterizing each filter, we put forward a hypernetwork (von Oswald et al., 2020) that adaptively generates filter parameters – MLP weights and biases in our case – for each target. When debiasing post embeddings with an identified target described by one or several keywords, the input to the hypernetwork is defined as the target’s indicator embedding, which can be easily composed with off-the-shelf pretrained word embeddings [^1] GetFair adopts GloVe (Pennington et al., 2014) word embeddings.. An adversarial training paradigm is in place to optimize the filters toward learning debiased post embeddings. Concretely, a target classifier is in place as the discriminator and tries to infer the original targets from the filtered post embeddings, where a capable filter function is ultimately learned to maximize the classification error of the discriminator. On the one hand, as the target embedding serves as the indicator to control the parameter generation process, the uniqueness of generated target-specific filters is guaranteed. On the other hand, by ensembling target-specific filters, GetFair is able to efficiently handle cases where a combination of different targets are identified in the same post, which is a trickier yet understudied case in real-world scenarios (Lalor et al., 2022). Moreover, the readily available target embeddings makes it possible for the hypernetwork to generate filter parameters with a valid indicator for an arbitrary, unseen target. To further refine the dynamic filter generation for each target that arrives, a semantic gap alignment scheme is proposed, such that the semantic distance among all targets’ input embeddings is resembled in the parameter space of their corresponding filters.
@@ 2. Preliminaries
Definition 1: Targets in A Post. In our setting, each social media post $s_{i}$ has one or more targeted user groups being discussed. In $s_{i}$, those mentioned user groups are termed targets, denoted by $t\in\mathcal{T}_{i}$. If deployed in a real online environment, one can consider using rule- or lexical-based approaches (Silva et al., 2016; ElSherief et al., 2018) or trainable detectors (Sharma et al., 2023) to identify the target set $\mathcal{T}_{i}$ from each post $s_{i}$. Because our emphasis is to uplift target-aware HSD fairness, we directly employ datasets (see Section 4.1 for details) that come with target labels, which are identified by human annotators from the textual clues within each post (Mathew et al., 2021; Ousidhoum et al., 2019; Sheth et al., 2024). The target identification task, as a more widely studied task, is not investigated in the paper.
Definition 2: Target-aware Fairness and Metrics. We assume a finite set of targeted groups $\mathcal{T}$ used for evaluation. For posts containing each target $t\in\mathcal{T}$, the overall group-level HSD classification performance can be measured by a specific accuracy metric. Then, fairness across targeted groups is reflected by the performance discrepancy among all $|\mathcal{T}|$ targets. Intuitively, the smaller the target-wise discrepancy, the fairer the HSD classifier. In our work, to quantify the level of algorithmic fairness, we adopt the well-established notion of False Positive/Negative Equality Difference (Dixon et al., 2018) suggested by Google Jigsaw, abbreviated as FPED/FNED. The only slight difference is that we normalize their values with the number of targets involved to make them comparable across different datasets, and the normalized versions are termed nFPED and nFNED:
@@
Definition 3: Hate Speech Detection with Generalizable Target-aware Fairness. Given a set of posts $\mathcal{S}_{train}=\{(s_{i},\mathcal{T}_{i},y_{i})\}_{i=1}^{|\mathcal{S}_{train}|}$ where the $i$-th post $s_{i}$ is associated with the binary label $y_{i}$ and a subset of known targets $\mathcal{T}_{i}\in\mathcal{T}_{train}$, $\mathcal{T}_{i}\neq\emptyset$, our objective is to train a debiased HSD classifier $\phi_{\text{HSD}}(\cdot)$. The trained classifier $\phi_{\text{HSD}}(\cdot)$ is evaluated on the test set $\mathcal{S}_{test}=\{(s_{j},\mathcal{T}_{j},y_{j})\}_{j=1}^{|\mathcal{S}_{test}|}$, where $\mathcal{T}_{j}\in\mathcal{T}_{test}$, $\mathcal{T}_{j}\neq\emptyset$. More importantly, all test-time posts will contain at least one target that is completely unseen during training, i.e., $\mathcal{T}_{test}\supset\mathcal{T}_{train}$. Ideally, $\phi_{\text{HSD}}(\cdot)$ is expected to: (1) provide maximal classification effectiveness, reflected by larger scores in accuracy and F1 score; and (2) yield minimal target-wise bias, reflected by lower scores in nFPED, nFNED, and HF.
@@ 3.1. Target-specific Filter Generation with Adaptive Hypernetwork
In GetFair, we utilize a filter function $f(\cdot)$ which takes the generated $d$-dimensional content embedding $\textbf{s}\in\mathbb{R}^{d}$ from an arbitrary encoder $g(\cdot)$ as its input, and emits a debiased representation $\tilde{\textbf{s}}=f(\textbf{s})\in\mathbb{R}^{d}$. The filter $f(\cdot)$ will be trained to remove target-specific information from the content embedding s, such that subsequent hate speech predictions made with $\tilde{\textbf{s}}$ is less reliant on it and will exhibit less performance bias on different targets while maintaining accuracy. Without loss of generality, $f(\cdot)$ can be formulated as a multi-layer perceptron (MLP) with weights $\textbf{W}\in\mathbb{R}^{d\times d}$ and biases $\textbf{b}\in\mathbb{R}^{d}$ at all layers. Considering the vast pool of possible targets in online platforms, it is unrealistic to rely on a single MLP’s capacity for removing all targets’ relevant information from s. Thus, a more performant approach is to instantiate one filter function $f_{t}(\cdot)$ per target $t$. However, this will inevitably lead to difficulties in scaling to numerous targets in reality and generalizing to unseen targets (e.g., an emerging buzzword) during inference.
@@
where $h_{l}(\cdot)$ is the hypernetwork responsible for generating parameters (i.e., weights and biases) for the $l$-th layer of filter $f_{t}(\cdot)$, and [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] denotes horizontal concatenation. For notation simplicity, we use $\Theta_{t}^{(l)}\in\mathbb{R}^{d\times(d+1)}$ to denote the concatenated weight and bias at each layer. Following common practices in hypernetworks (Yan et al., 2022; Shamsian et al., 2021), $h_{l}(\cdot)$ firstly outputs a flat vector with $d\times(d+1)$ dimensionality, which is further reshaped into the matrix form of $\Theta_{t}^{(l)}$.
Target Indicators. Notably, $h_{l}(\cdot)$ is shared across all targets, and the generated parameters are solely conditioned on its input $\textbf{t}\in\mathbb{R}^{d^{\prime}}$, which is the target indicator that informs the hypernetwork of the exact target filter to generate. To facilitate effective information filtering and target-wise generalizability, the design of the target indicators should meet two characteristics: (1) they should be distinguishable for different targets such that every target-specific filter maintains uniqueness; and (2) they are capable of representing an arbitrary number of unseen targets without any training. Hence, this eliminates some indicator methods commonly used in hypernetworks such as one-hot encoding and learnable embeddings (von Oswald et al., 2020; Yan et al., 2022; Shamsian et al., 2021). In GetFair, we take advantage of the wide availability of pretrained word embeddings, namely the GloVe embeddings (Pennington et al., 2014) for composing the target indicator t. We have adopted its $300$-dimensional version, thus $d^{\prime}=300$. Specifically, for target $t=\{v_{1},...,v_{k},...,v_{n}\}$ represented as a sequence of $n$ vocabularies, we take the mean of all $n$ corresponding word embeddings $\textbf{v}_{k}\in\mathbb{R}^{d^{\prime}}$ as its representation, i.e., $\textbf{t}=\frac{1}{n}\sum_{k=1}^{n}\textbf{v}_{k}$.
@@
Multi-target Filters. It is worth mentioning that, in real applications, one post might be attributed to several different targets, calling upon the necessity for a mechanism that can simultaneously filter a set of targets $\mathcal{T}_{i}$ associated with each post’s embedding $\textbf{s}_{i}$. A straightforward way of dosing so is to remove one target-specific information at a time via $\tilde{\textbf{s}}_{i,t}=f_{t}(\textbf{s}_{i})$ for every $t\in\mathcal{T}_{i}$, then merge all filtered embeddings of the same post (e.g., via sum pooling (Li et al., 2021)) into $\tilde{\textbf{s}}_{i}$. Though each $\tilde{\textbf{s}}_{i,t}$ removes information related to target $t\in\mathcal{T}_{i}$, it is not necessarily the case for a different $\tilde{\textbf{s}}_{i,t^{\prime}}$ where $t^{\prime}\in\mathcal{T}_{i}$ and $t^{\prime}\neq t$, as the associated filter $f_{t^{\prime}}(\cdot)$ is only dedicated to removing information about $t^{\prime}$. Consequently, the fused embedding fusion step may introduce unwanted target-specific information back into the final embedding $\tilde{\textbf{s}}_{i}$. To bypass this potential defect, we design a combinatorial approach to form a multi-target filter via parameter ensemble:
@@ 3.2. Regularizing Filter Parameters with Semantic Gap Alignment
Compared with learning multiple filter functions, learning a unified adaptive hypernetwork brings substantially lighter parameterization, but also comes at a risk of overfitting some specific target distributions, e.g., the frequent ones, challenging the generlizability. Hence, the hypernetwork, if trained without regularization, can potentially fail to generalize when generating filter weights for completely unseen targets. To prevent the generated filter parameters from undesired homogeneity and thus impaired utility/generalizability, we impose an additional regularization on the distribution of the generated filter parameters $\Theta_{t}^{(l)}$ w.r.t. different targets. Specifically, for every pair of targets $t$, $t^{\prime}$ and their indicators t, $\textbf{t}^{\prime}$, we would like to preserve their semantic distance within the generated filter parameters parameters $\Theta_{t}$, $\Theta_{t^{\prime}}$ too, i.e., $d(\textbf{t},\textbf{t}^{\prime})\approx d(\Theta_{t}^{(l)},\Theta_{t^{\prime}}^{(l)})$ with a distance metric $d(\cdot,\cdot)$. To bypass the unmatched dimensionality between t and $\Theta_{t}$, we define a semantic gap alignment scheme as the regularization loss $\mathcal{L}_{reg}$ as follows:
@@ 3.3. Target Discriminator
Ideally, each debiased post embedding $\tilde{\textbf{s}}_{i}$ no longer carries target-specific information that leads to the hypothetically unfair classification results. To ensure that the target filter removes the targeted-related information, we propose to take advantage of adversarial learning to optimize each filter. Specifically, we set up a target discriminator defined as the following:

### v2

@@ Abstract.
To counter the side effect brought by the proliferation of social media platforms, hate speech detection (HSD) plays a vital role in halting the dissemination of toxic online posts at an early stage. However, given the ubiquitous topical communities on social media, a trained HSD classifier can easily become biased towards specific targeted groups (e.g., female and black people), where a high rate of either false positive or false negative results can significantly impair public trust in the fairness of content moderation mechanisms, and eventually harm the diversity of online society. Although existing fairness-aware HSD methods can smooth out some discrepancies across targeted groups, they are mostly specific to a narrow selection of targets that are assumed to be known and fixed. This inevitably prevents those methods from generalizing to real-world use cases where new targeted groups constantly emerge (e.g., new forums created on Reddit) over time. To tackle the defects of existing HSD practices, we propose Generalizable target-aware Fairness (GetFair), a new method for fairly classifying each post that contains diverse and even unseen targets during inference. To remove the HSD classifier’s spurious dependence on target-related features, GetFair trains a series of filter functions in an adversarial pipeline, so as to deceive the discriminator that recovers the targeted group from filtered post embeddings. To maintain scalability and generalizability, we innovatively parameterize all filter functions via a hypernetwork. Taking a target’s pretrained word embedding as input, the hypernetwork generates the weights used by each target-specific filter on-the-fly without storing dedicated filter parameters. In addition, a novel semantic gap alignment scheme is imposed on the generation process, such that the produced filter function for an unseen target is rectified by its semantic affinity with existing targets used for training. Finally, experiments [^1] The implementation of GetFair is released at https://github.com/xurong-liang/GetFair. are conducted on two benchmark HSD datasets, showing advantageous performance of GetFair on out-of-sample targets among baselines.
@@
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 25–29, 2024; Barcelona, Spainbooktitle: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining (KDD ’24), August 25–29, 2024, Barcelona, Spaindoi: 10.1145/3637528.3671821isbn: 979-8-4007-0490-1/24/08ccs: Information systems Web miningccs: Information systems Data miningccs: Security and privacy Social aspects of security and privacy
@@ 1. Introduction
To this end, in this paper we propose a novel approach, namely HSD with Generalizable target-aware Fairness (GetFair). In GetFair, to avoid the linear parameter cost associated with the number of target-specific filter functions, instead of independently parameterizing each filter, we put forward a hypernetwork (von Oswald et al., 2020) that adaptively generates filter parameters – MLP weights and biases in our case – for each target. When debiasing post embeddings with an identified target described by one or several keywords, the input to the hypernetwork is defined as the target’s indicator embedding, which can be easily composed with off-the-shelf pretrained word embeddings [^2] GetFair adopts GloVe (Pennington et al., 2014) word embeddings.. On the one hand, as the target embedding serves as the indicator to control the parameter generation process, the uniqueness of generated target-specific filters is guaranteed. On the other hand, by ensembling target-specific filters, GetFair is able to efficiently handle cases where a combination of different targets are identified in the same post, which is a trickier yet understudied case in real-world scenarios (Lalor et al., 2022). Moreover, the readily available target embeddings makes it possible for the hypernetwork to generate filter parameters with a valid indicator for an arbitrary, unseen target. An adversarial training paradigm is in place to optimize the filters toward learning debiased post embeddings. Concretely, a target classifier is deployed as the discriminator and tries to infer the original targets from the filtered post embeddings, where a capable filter function is ultimately learned to maximize the classification error of the discriminator. To further refine the dynamic filter generation for each target that arrives, a semantic gap alignment scheme is proposed, such that the semantic distance among all targets’ input embeddings is resembled in the parameter space of their corresponding filters.
@@ 2. Preliminaries
Definition 1: Targets in A Post. In our setting, each social media post $s_{i}$ has one or more targeted user groups being discussed. In $s_{i}$, those mentioned user groups are termed targets, denoted by $t\in\mathcal{T}_{i}$. If deployed in a real online environment, one can consider using rule- or lexical-based approaches (Silva et al., 2016; ElSherief et al., 2018), or trainable detectors (Sharma et al., 2023) to identify the target set $\mathcal{T}_{i}$ from each post $s_{i}$. Because our emphasis is to uplift target-aware HSD fairness, we directly employ datasets (see Section 4.1 for details) that come with target labels, which are identified by human annotators from the textual clues within each post (Mathew et al., 2021; Ousidhoum et al., 2019; Sheth et al., 2024). The target identification task, as a more widely studied task, is not investigated in the paper.
Definition 2: Target-aware Fairness Metrics. We assume a finite set of targeted groups $\mathcal{T}$ used for evaluation. For posts containing each target $t\in\mathcal{T}$, the overall group-level HSD classification performance can be measured by a specific accuracy metric. Then, fairness across targeted groups is reflected by the performance discrepancy among all $|\mathcal{T}|$ targets. Intuitively, the smaller the target-wise discrepancy, the fairer the HSD classifier. In our work, to quantify the level of algorithmic fairness, we adopt the well-established notion of False Positive/Negative Equality Difference (Dixon et al., 2018) suggested by Google Jigsaw, abbreviated as FPED/FNED. The only slight difference is that we normalize their values with the number of targets involved to make them comparable across different datasets, and the normalized versions are termed nFPED and nFNED:
@@
Definition 3: Hate Speech Detection with Generalizable Target-aware Fairness. Given a set of posts $\mathcal{S}_{train}=\{(s_{i},\mathcal{T}_{i},y_{i})\}_{i=1}^{|\mathcal{S}_{train}|}$ where the $i$-th post $s_{i}$ is associated with the binary label $y_{i}$ and a subset of known targets $\mathcal{T}_{i}\in\mathcal{T}_{train}$, $\mathcal{T}_{i}\neq\emptyset$, our objective is to train a debiased HSD classifier $\phi_{\text{HSD}}(\cdot)$. The trained classifier $\phi_{\text{HSD}}(\cdot)$ is evaluated on the test set $\mathcal{S}_{test}=\{(s_{j},\mathcal{T}_{j},y_{j})\}_{j=1}^{|\mathcal{S}_{test}|}$, where $\mathcal{T}_{j}\in\mathcal{T}_{test}$, $\mathcal{T}_{j}\neq\emptyset$. More importantly, all test-time posts will contain at least one target that is completely unseen during training, i.e., $\mathcal{T}_{test}\setminus\mathcal{T}_{train}\neq\emptyset$. Ideally, $\phi_{\text{HSD}}(\cdot)$ is expected to: (1) provide maximal classification effectiveness, reflected by larger scores in accuracy, F1, and AUC; and (2) yield minimal target-wise bias, reflected by lower scores in nFPED, nFNED, and HF.
@@ 3.1. Target-specific Filter Generation with Adaptive Hypernetwork
In GetFair, we utilize a filter function $f(\cdot)$ which takes the generated $d$-dimensional content embedding $\textbf{s}\in\mathbb{R}^{d}$ from an arbitrary encoder $g(\cdot)$ as its input, and emits a debiased representation $\tilde{\textbf{s}}=f(\textbf{s})\in\mathbb{R}^{d}$. The filter $f(\cdot)$ will be trained to remove target-specific information from the content embedding s, such that subsequent hate speech predictions made with $\tilde{\textbf{s}}$ are less reliant on it and will exhibit less performance bias on different targets while maintaining accuracy. Without loss of generality, $f(\cdot)$ can be formulated as a multi-layer perceptron (MLP) with weights $\textbf{W}\in\mathbb{R}^{d\times d}$ and biases $\textbf{b}\in\mathbb{R}^{d}$ at all layers. Considering the vast pool of possible targets in online platforms, it is unrealistic to rely on a single MLP’s capacity for removing all targets’ relevant information from s. Thus, a more performant approach is to instantiate one filter function $f_{t}(\cdot)$ per target $t$. However, this will inevitably lead to difficulties in scaling to numerous targets in reality and generalizing to unseen targets (e.g., an emerging buzzword) during inference.
@@
where $h_{l}(\cdot)$ is the hypernetwork responsible for generating parameters (i.e., weights and biases) for the $l$-th layer of filter $f_{t}(\cdot)$, and [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] denotes horizontal concatenation. For notation simplicity, we use $\Theta_{t}^{(l)}\in\mathbb{R}^{d\times(d+1)}$ to denote the concatenated weight and bias at each layer. Following common practices in hypernetworks (Yan et al., 2022; Shamsian et al., 2021), $h_{l}(\cdot)$ firstly outputs a flat vector with $d^{2}+d$ dimensionality, which is further reshaped into the matrix form of $\Theta_{t}^{(l)}$.
Target Indicators. Notably, $h_{l}(\cdot)$ is shared across all targets, and the generated parameters are solely conditioned on its input $\textbf{t}\in\mathbb{R}^{d^{\prime}}$, which is the target indicator that informs the hypernetwork of the exact target filter to generate. To facilitate effective information filtering and target-wise generalizability, the designed target indicators should meet two characteristics: (1) they should be distinguishable for different targets such that every target-specific filter maintains uniqueness; and (2) they are capable of representing an arbitrary number of unseen targets without any training. Hence, this eliminates some indicator methods commonly used in hypernetworks such as one-hot encoding and learnable embeddings (von Oswald et al., 2020; Yan et al., 2022; Shamsian et al., 2021). In GetFair, we take advantage of the wide availability of pretrained word embeddings, namely the GloVe embeddings (Pennington et al., 2014) for composing the target indicator t. We have adopted its $300$-dimensional version, thus $d^{\prime}=300$. Specifically, for target $t=\{v_{1},...,v_{k},...,v_{n}\}$ represented as a sequence of $n$ vocabularies, we take the mean of all $n$ corresponding word embeddings $\textbf{v}_{k}\in\mathbb{R}^{d^{\prime}}$ as its representation, i.e., $\textbf{t}=\frac{1}{n}\sum_{k=1}^{n}\textbf{v}_{k}$.
@@
Multi-target Filters. It is worth mentioning that, in real applications, one post might be attributed to several different targets, calling upon the necessity for a mechanism that can simultaneously filter a set of targets $\mathcal{T}_{i}$ associated with each post’s embedding $\textbf{s}_{i}$. A straightforward way of dosing so is to remove one target-specific information at a time via $\tilde{\textbf{s}}_{i,t}=f_{t}(\textbf{s}_{i})$ for every $t\in\mathcal{T}_{i}$, then merge all filtered embeddings of the same post (e.g., via sum pooling (Li et al., 2021)) into $\tilde{\textbf{s}}_{i}$. Though each $\tilde{\textbf{s}}_{i,t}$ removes information related to target $t\in\mathcal{T}_{i}$, it is not necessarily the case for a different $\tilde{\textbf{s}}_{i,t^{\prime}}$ where $t^{\prime}\in\mathcal{T}_{i}$ and $t^{\prime}\neq t$, as the associated filter $f_{t^{\prime}}(\cdot)$ is only dedicated to removing information about $t^{\prime}$. Consequently, the embedding fusion step may introduce unwanted target-specific information back into $\tilde{\textbf{s}}_{i}$. To bypass this potential defect, we design a combinatorial approach to form a multi-target filter via parameter ensemble:
@@ 3.2. Regularizing Filter Parameters with Semantic Gap Alignment
Compared with learning multiple filter functions, learning a unified adaptive hypernetwork brings substantially lighter parameterization, but also comes at a risk of overfitting some specific target distributions, e.g., the frequent ones, challenging the generlizability. Hence, the hypernetwork, if trained without regularization, can potentially fail to generalize when generating filter weights for completely unseen targets. To prevent the generated filter parameters from undesired homogeneity and thus impaired utility, we impose an additional regularization on the distribution of the generated filter parameters $\Theta_{t}^{(l)}$ w.r.t. different targets. Specifically, for every pair of targets $t$, $t^{\prime}$ and their indicators t, $\textbf{t}^{\prime}$, we would like to preserve their semantic distance within the generated filter parameters $\Theta_{t}$, $\Theta_{t^{\prime}}$ too, i.e., $d(\textbf{t},\textbf{t}^{\prime})\approx d(\Theta_{t}^{(l)},\Theta_{t^{\prime}}^{(l)})$ with a distance metric $d(\cdot,\cdot)$. To bypass the unmatched dimensionality between t and $\Theta_{t}$, we define a semantic gap alignment scheme as the regularization loss $\mathcal{L}_{reg}$ as follows:
@@ 3.3. Target Discriminator
Ideally, each debiased post embedding $\tilde{\textbf{s}}_{i}$ no longer carries target-specific information that leads to the hypothetically unfair classification results. To ensure that the target filter removes the target-related information, we propose to take advantage of adversarial training to optimize each filter. Specifically, we set up a target discriminator defined as the following: